OpenStudy (anonymous):

If f(1/x) = 2x, what is f(x)? (a) 2/x (b) x/2 (c) 1/2x (d) 2 (e) cannot be determined

OpenStudy (australopithecus):

f(1/x) = 2x then we have f(x) = 2/x Lets test it if you plug 1/x into f(x) = 2/(1/x) you get f(1/x) = 2/(1/x) = (2/1)(x/1) = 2x

OpenStudy (australopithecus):

just use logic to solve these problems

OpenStudy (australopithecus):

problem solved

OpenStudy (australopithecus):

Lets test it if you plug 1/x into f(x) = 2/x you get f(1/x) = 2/(1/x) = (2/1)(x/1) = 2x

OpenStudy (anonymous):

thank you :)